OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS

OpenStudy (anonymous):

thats all it gives

OpenStudy (anonymous):

i'm assuming no REAL solutions.... \[\large \sqrt{x}=-2 \]

OpenStudy (anonymous):

Ok, how about you write up a quadratic equation with no real solutions.

OpenStudy (anonymous):

idk what to do..

OpenStudy (anonymous):

Do what dpalnc said.

OpenStudy (anonymous):

\[10x^2-x+20\] Notice using the quadratic equation gives a negative square root value? Which is what @dpaInc just showed you.

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

Bryan, do this one then. x-1 = x + 1

OpenStudy (anonymous):

It results in - 1 = 1 which is not true and thus no solutions.

OpenStudy (anonymous):

oh no wonder i was like i cant solve that

OpenStudy (anonymous):

thats not two step though or is it?

OpenStudy (anonymous):

Medal please? Lol

OpenStudy (anonymous):

is taht a two step equation???

OpenStudy (anonymous):

3(2x + 5) = 2(3x + 4) Ok, this is a two step equation.

OpenStudy (anonymous):

Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (anonymous):

3(2x+5) = 2(3x+4) 6x + 15 = 6x + 8 15 = 8 This proves there are no solutions because the end result is 15 = 8, which is clearly not true. For example, this is saying 15 cats are equal to 8 cats -- obviously a false statement.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Medal please.